A toy car slides straight off a 4.1 m high book shelf at a velocity of 2.8 m/s.

a) How long does it take for the toy car to hit the ground?

b) what is the range of the car?


1
Expert's answer
2021-12-03T12:45:08-0500

Given:

"v_0=2.8\:\rm m/s"

"H=4.1\:\rm m"


a) the time of motion

"t=\sqrt{2H/g}=\rm\sqrt{2*4.1\: m/9.8\: m/s^2}=0.91\: s"

b) the range

"L=v_0t=\rm 2.8\: m/s*0.91\: s=2.6\: m"